Divide
(x^2+11x+18) divided by (x+2)

Found 2 solutions by richwmiller, xgeumjandi:
Answer by richwmiller(17219)

Since (x^2+11x+18) =(x+9)* (x+2)
(x^2+11x+18) divided by (x+2)=x+9

Answer by xgeumjandi(1)

x^2 + 11x + 18 divde by x+2 (x+2 becomes -2)
1.take the first number out of each term:
x^2 =1, 11x=11, 18= 18
2.divide each term by -2
-2) 1 11 18
-2 -18
_________________
1 9 0
x^2 = turns to x
11x = turns to 9
18= turns to 0
the answer wil be: x+9